OpenStudy (anonymous):

Write an equation for the following table: x=time(s) y=costs x: 0,5,10,15,30,40,40,50,50,60 y: 38,51,46,73,60,79,93,77,100,100 a. What will the costs be if the time is 22 minutes? Please use the equation

OpenStudy (anonymous):

help please!

OpenStudy (anonymous):

Which equation are you directed to use?

OpenStudy (anonymous):

hummm..i just need the equation for he table!

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

Sure, it could save me a bit of time if you tell me what you're working on in class

OpenStudy (anonymous):

Well, for this question..i need to find the y-intercept and slope and write the equation! But what i work in class right now is a review of everything...like, probability, slope, y-intercept,inequality,equation, else...

OpenStudy (anonymous):

Did you copy all of the numbers correctly? That is not a linear equation, so unless you need the derivative, it will be hard (impossible) to determine a 'slope'.

OpenStudy (anonymous):

And the derivative is also kind or meaningless.

OpenStudy (anonymous):

really!?

OpenStudy (anonymous):

Yes. It would be possible to create an equation for the average change between x=20 and x=25, but that may not be what your teacher wants.

OpenStudy (anonymous):

Or, 15 and 30..

OpenStudy (anonymous):

If that is what we are solving for, we could just plug numbers into the point-slope form of the equation for a line.

OpenStudy (anonymous):

Namely, (y - y1) = m (x - x1) Where m is the slope of the line (in this case, from (15,73) to (30,60)), and y1 and x1 are the y and x coordinates of one of those points. If you put it into the form, y=mx+b, you will have the slope and y intercept. If you plug in x=22, you will have your answer.

OpenStudy (anonymous):

Do you understand that?

OpenStudy (anonymous):

i understand that but i also agree with you that it could not write the equation. i forgot that my teacher also said that "graph the following table and draw a trend line then write an equation for it"

OpenStudy (anonymous):

Do you truly find my prose to be of such dismal quality?

OpenStudy (anonymous):

Ah, well, this dataset kind of sucks for drawing a good trend line, but as long as you remember what they want, good.

OpenStudy (anonymous):

yupp. well. thank you so much anyway :)

OpenStudy (anonymous):

Sure, no problem